if you react 120g of c2h6 with 320g of o2 what is the limiting reagent?


1
Expert's answer
2022-02-02T02:15:02-0500

Molar mass of "O_2=32g"

Moles of "O_2=\frac{320}{32}=10"


Molar mass of "C_2H_6=30g"

Moles of "C_2H_6=\frac{120}{30}=4"


"2C_2H_6+7O_2\to 4CO_2+6H_2O"


"4" moles of "C_2H_6" requires

"\frac {4}{2}×7=14" moles of "O_2"


The available moles of "O_2" are 10


"10" moles of "O_2" requires

"\frac{10}{7}×2= 2.8" moles of "C_2H_6"


Therefore, "O_2" is the limiting reagent.